Management of a firm with a cost of capital of 10 percent is considering a $126,000 investment with annual cash flow of $52,460 for three years. Use Appendix A and Appendix D to answer the questions.

What are the investment’s net present value and internal rate of return? Use a minus sign to enter a negative value, if any. Round your answers for the net present value to the nearest dollar and for the internal rate of return to the nearest whole number.

NPV: = $(?)

IRR: = %(?)

The internal rate of return assumes that each cash flow is reinvested at the internal rate of return. If that reinvestment rate is achieved, what is the total value of the cash flows at the end of the third year? Use the rounded internal rate of return from part a. Round your answer to the nearest dollar.

$ = (?)

The net present value technique assumes that each cash flow is reinvested at the firm’s cost of capital. What would be the total value of the cash flows at the end of the third year, if the funds are reinvested at the firm’s cost of capital? Round your answer to the nearest dollar. $ = (?)
Answer #1

Below is the calculation of NPV and IRR

Present Value of Cash 1 Year Cash Flow Flow(Cash Flow/1.1^year) 0 (126,000.00) (126,000.00) 1 52,460.00 47,690.91 2 52,460.00

Awn- B. Present Value of Cash 1 Year Cash Flow Flow(Cash Flow/1.1^year) (126,000.00) (126,000.00) 52,460.00 47,690.91 52,460.

Hence, NPV= $4460.26 and IRR=12%

If investment rate is IRR i.e. 12%, then value of cash flow after 3 year= 52460*(1+12%)^2+52460*(1+12%)+52460=$177021.02

Similarly, if the cash flow are reinvested at firm's cost of capital i.e. 10%, then value of cash flows at the end of third year= 52460*(1+10%)^2+52460*(1+10%)+52460=$173642.60